OpenStudy (lena772):

The first four terms of a sequence are shown below. 7, 4, 1, -2 Which of the following functions best defines this sequence? f(1) = 7, f(n + 1) = f(n) + 3; for n ≥ 1 f(1) = 7, f(n + 1) = f(n) - 3; for n ≥ 1 f(1) = 7, f(n + 1) = f(n) - 4; for n ≥ 1 f(1) = 7, f(n + 1) = f(n) + 4; for n ≥ 1

OpenStudy (the_fizicx99):

B?

OpenStudy (lena772):

eep idk

OpenStudy (lena772):

@AllTehMaffs

OpenStudy (the_fizicx99):

f(1+1) = f(2) - 3 = 1 f(2+1) = f(3) - 3 = -2

OpenStudy (the_fizicx99):

I think it should've been f(0+1) = f(1) which is 7 - 3 which is 4.

OpenStudy (the_fizicx99):

@Lena772 you get it? 4 is f(2) - 3 is 1 f(3) is 1 - 3 is -2

OpenStudy (lena772):

no :/

OpenStudy (anonymous):